OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet.

OpenStudy (anonymous):

How many bracelets were sold?

OpenStudy (anonymous):

8 bracelets were sold.

OpenStudy (anonymous):

54-6=48 so the bracelets by them self costed $48 48/8=6 so each bracelet was $6

OpenStudy (goldphenoix):

You can use an equation to solve this. \[\large \large 8x + 6 = 54\]